11)A supply curve that is horizontally sloped is an accurate representation of the law of supply given ceteris paribus when only price changes for a good is a:

(a)False statement

(b)True statement

(c)More information needed to respond

(d)All of the above

(12)Positive economic analysis deals with_________________________ and normative economic analysis deals with_____________________

(a)What should be, what is

(b)What is, what ought to be

(c)Fiction, fact

(d)Positive things, negative things

(13)If given ceteris paribus there is a positive correlation between annual household expenditures and annual household incomes in an economy, macroeconomists will conclude that:

(a)A decrease in household incomes will cause a decrease in household consumption expenditures and vice versa

(b)An increase in household incomes will cause a decrease in household consumption expenditures and vice versa

(c)A decrease in household incomes will have no effect on household consumption expenditures and vice versa

(d)None of the above

(14)Suppose in a market for organic carrots, you are asked to compute the market clearing price and market clearing quantity for airline tickets in a market where the demand and supply equations are D=340-5P and S=250, which of the following pairs would be your accurate calculations of what you are asked to do?

(a)P=$18 and Q=156 tickets

(b)P=$12 and Q=280 tickets

(c)P=$9 and Q=250 tickets

(d)None of the above

(15)Suppose now, there is a change in the market conditions in the market for airline tickets in Q#14 above such that the supply of tickets is represented by the equation S=120 and the demand for tickets is unchanged as described in the previous question, which of the following is the market equilibrium price and the market equilibrium quantity in the market?

(a)P=$15 and Q=120 tickets

(b)P=$20 and Q=280 tickets

(c)P=$44 and Q=120 tickets

(d)P=$35 and Q=320 tickets

(16)Which of the following pairs of statements is true about the slopes of the demand and supply “curves” in Qs#14 and 15 above?

(a)The demand curve is positively sloped and the supply curve is negatively sloped

(b)The demand curve is negatively sloped and the supply curve is vertically sloped

(c)The demand curve is negatively sloped and the supply curve is horizontally sloped

(d)(b) and (c) above

(17)The silly clothes worn by circus clowns are considered as an example of:

(a)Human resources

(b)Natural resources

(c)Capital resources

(d)Entrepreneurship

(18)Which of the following statements is true?

Market prices are those prices set by the market forces of supply and demand in any market:

(a)At equilibrium

(b)Above equilibrium

(c)Below equilibrium

(d)All of the above

(19)Which of the following statements is true?

If there is a direct relationship between the price level and the money supply in an economy, it would be correct to conclude that:

(a)There is an inverse relationship between the price level and the money supply in the economy

(b)There is a vertical relationship between the price level and the money supply in the economy

(c)A decrease in the price level leads to an increase in the money supply in the economy

(d)None of the above

(20)Which of the following statements is false?

(a)A rent control is an example of a binding price ceiling

(b)A binding price ceiling is set above a competitive equilibrium price

(c)A legally binding minimum wage is set above the equilibrium wage rate in a labor market

(d)None of the above
Answer #1

11. The law of supply states that, ceteris paribus, an increase in price results in an increase in quantity supplied.

a) False statement.

12. b) what is, what ought to be.

13. a) a decrease in househilhincomes will cause a decrease in household consumption expenditure and vice versa.

14. D= 340-5P

S = 250.

At equilibrium, demand = supply.

Therefore, 340-5P = 250

90= 5P

P=90/5 = $ 18

Equilibrium price is $18 and equilibrium quantity is 250 units.

d) none of the above.